OpenStudy (anonymous):

What is the equation of the line passing through (0, 2) and perpendicular to the line y = 4x – 3 in slope-intercept form?

OpenStudy (anonymous):

If it is perpendicular to that line the perpendicular line has a slope that is a negative reciprocal to 4 meaning -1/4 If you substitute into y = mx + b 2 = (-1/4)(0) + b so b = 2 equation is y = -1/4 x + 2